mat 115 chapter 4 in basic mathematical skills with geometry
Business and finance. Claudia earns $6.40 per hour (h). For overtime (each hour
over 40 h) she earns $9.60. If she works 48.5 h in a week, what pay should she
receive?

Answer by checkley77(12844)

40*6.40+8.5*9.60=256+81.60=$337.60 IS HER TOTAL PAY.